### v1

# Gamma family characterization and an alternative proof of Gini estimator unbiasedness

Abstract
In this paper, we derive a general representation for the expectation of the (upward adjusted) Gini coefficient estimator in terms of the Laplace transform of the underlying distribution. This representation leads to a characterization of the gamma family within the class of nonnegative scale families, based on a stability property of exponentially tilted distributions. As an application, we provide an alternative proof of the unbiasedness of the Gini coefficient estimator under gamma populations.
Keywords. Gamma distribution, Gini coefficeint, Gini estimator, unbiased estimator.
Mathematics Subject Classification (2010). MSC 60E05 $\cdot$ MSC 62Exx $\cdot$ MSC 62Fxx.
1 Introduction
The Gini coefficient is a widely used measure of inequality in economics, finance, demography, and reliability theory, and has been extensively studied since its introduction by Gini, (1936), particularly with regard to the statistical properties of its estimators.
A well-known issue is the finite-sample bias of the classical Gini estimator. To address it, Deltas, (2003) proposed an upward-adjusted estimator that corrects small-sample bias under particular distributional assumptions. Its unbiasedness has since been established for several distribution families, most notably the gamma family; see Baydil et al., (2025) and Vila and Saulo, 2025a ; Vila and Saulo, 2025b .
This paper makes two complementary contributions. First, we derive a general expression for the expectation of the Gini estimator for any nonnegative, absolutely continuous distribution with finite mean. This approach naturally connects to the Gini mean difference (GMD) studied in Vila et al., (2024), highlighting how the distributional properties of the underlying population influence the Gini-based measures of variability and inequality. Second, we show that a natural scale-family invariance condition on the associated exponentially tilted distribution characterizes the gamma family.
This characterization yields a unified proof of the unbiasedness of the Gini estimator under gamma populations and clarifies the fundamental role of scale invariance in explaining the prominence of the gamma family in Gini estimation.
The remainder of the paper is organized as follows. In Section 2 we review basic properties and representations of the Gini coefficient. In Section 3 we derive a general expression for the expectation of the Gini estimator. In Section 4 we establish a characterization of the gamma family within the class of scale families. In Section 5 we apply these results to obtain an alternative proof of the unbiasedness of the Gini estimator under gamma populations. Finally, in Section 6, we provide some concluding remarks.
2 The Gini coefficient
Let $X_{1},X_{2}$ be independent and identically distributed (i.i.d.) non-negative random variables with common distribution as a nonnegative, absolutely continuous random variable $X$, and finite mean $\mu=\mathbb{E}[X]>0$. The Gini coefficient (Gini,, 1936) is defined as
$G\equiv G(X)=\frac{\mathbb{E}\lvert X_{1}-X_{2}\rvert}{2\mu}.$
The Gini coefficient possesses the following well-known fundamental properties and representations:
  • Existence: the Gini coefficient exists and satisfies $0\leqslant G<1$.
  • Scale invariance: for any $b>0$, $G(bX)=G(X).$
  • Lack of translation invariance: for any $a>0$, $G(a+X)={\mu}G(X)/(a+\mu).$
  • Covariance and quantile representations: $G=\mathrm{Cov}\!\left(X,\,2F_{X}(X)-1\right)/\mu=\int_{0}^{1}F_{X}^{-1}(p)(2p-1)\,\mathrm{d}p,$ where $F_{X}$ is the cumulative distribution function (CDF) of $X$ and $F_{X}^{-1}$ its generalized inverse function.
  • Decomposition via Lorenz measures of inequality: $G=[D_{1}(F_{X})+G_{2}(F_{X})]/2,$ where $D_{n}(F_{X})=(n+1)\mathbb{E}\left[(U-L(U))U^{\,n-1}\right]$ and $G_{n}(F_{X})=n(n-1)\mathbb{E}\left[(U-L(U))(1-U)^{\,n-2}\right]$, $n\geqslant 1$, are the Lorenz measure of inequality (Aaberge,, 2000) and the generalized Gini measure (Donaldson and Weymark,, 1980; Kakwani,, 1980; Yitzhaki,, 1983), respectively, with $U\sim\mathrm{U}(0,1)$, and $L(p)=\int_{0}^{p}F_{X}^{-1}(t)\,\mathrm{d}t/\mu,\,0\leqslant p\leqslant 1,$ being the Lorenz curve associated with $X$.
The following result is well known in the literature; for completeness and the reader’s convenience, we briefly outline the main idea of its proof.
Proposition 2.1.
The Gini coefficient $G$ of a nonnegative, absolutely continuous random variable $X$ can be characterized as
$G={\int_{0}^{\infty}F_{X}(x)[1-F_{X}(x)]{\rm d}x\over\mu},$
where $\mu=\mathbb{E}[X]>0$ is the mean of $X$, and $F_{X}$ is the CDF of $X$.
Proof.
The proof is immediate upon using the basic identity:
$|x-y|=\int_{0}^{\infty}\left[1-\mathds{1}_{\{x\leqslant t,y\leqslant t\}}-\mathds{1}_{\{x\geqslant t,y\geqslant t\}}\right]{\rm d}t.$ (2.1)
3 A simple expression for the expectation of the Gini estimator
In this section (Theorem 3.1), we derive a general expression for the expectation of the (upward adjusted) Gini coefficient estimator $\widehat{G}$, initially proposed by Deltas, (2003).
Let $X_{1},\ldots,X_{n}$ be i.i.d. copies of $X$. The Gini coefficient estimator of $G$ is defined by
$\widehat{G}=\dfrac{\displaystyle{1\over\binom{n}{2}}\sum_{1\leqslant i<j\leqslant n}|X_{i}-X_{j}|}{2\overline{X}},$
where $\overline{X}=\sum_{k=1}^{n}X_{k}/n$ denotes the sample mean.
Theorem 3.1.
Let $X_{1},\ldots,X_{n}$ be independent copies of a nonnegative, absolutely continuous random variable $X$ with finite mean $\mu=\mathbb{E}[X]>0$ and CDF $F_{X}$. Then
$\mathbb{E}\left[\widehat{G}\right]=n\int_{0}^{\infty}\mathbb{E}[Y_{z}]\,G(Y_{z})\,\mathscr{L}_{X}^{\,n}(z)\,\mathrm{d}z,$
where $\mathscr{L}_{X}(z)=\mathbb{E}\left[\exp\{-zX\}\right]$ denotes the Laplace transform of $X$. For each $z>0$, $G(Y_{z})$ is the Gini coefficient of the random variable $Y_{z}$, whose CDF is given by
$F_{Y_{z}}(t)=\frac{\mathbb{E}\left[\exp\{-zX\}\mathds{1}_{\{X\leqslant t\}}\right]}{\mathscr{L}_{X}(z)},\quad t>0.$ (3.1)
In the above, we implicitly assume that the Lebesgue–Stieltjes and improper integrals involved exist.
Proof.
Using the identity $\int_{0}^{\infty}\exp(-wz){\rm d}z={1/w},\,w>0,$ together with Tonelli’s theorem, we can express $\mathbb{E}\left[\widehat{G}\right$] as follows
$\mathbb{E}\left[\widehat{G}\right$] $={n\over 2\binom{n}{2}}\sum_{1\leqslant i<j\leqslant n}\int_{0}^{\infty}\mathbb{E}\left[|X_{i}-X_{j}|\exp\{-z(X_{1}+X_{2})\}\exp\left\{-z\sum_{k=3}^{n}X_{k}\right\}\right]{\rm d}z$
$={n\over 2}\int_{0}^{\infty}\mathbb{E}\left[|X_{1}-X_{2}|\exp\{-z(X_{1}+X_{2})\}\right]\mathscr{L}_{F}^{n-2}(z){\rm d}z,$
where, in the last equality, we used the fact that $X_{1},\ldots,X_{n}$ are i.i.d. copies of $X$. Applying the identity in (2.1), the above identity can be written as
$\mathbb{E}\left[\widehat{G}\right]={n\over 2}\int_{0}^{\infty}\left\{\int_{0}^{\infty}\left\{\mathscr{L}_{F}^{2}(z)-\mathbb{E}^{2}\left[\exp\{-zX\}\mathds{1}_{\{X\leqslant t\}}\right]-\mathbb{E}^{2}\left[\exp\{-zX\}\mathds{1}_{\{X\geqslant t\}}\right]\right\}{\rm d}t\,\mathscr{L}_{F}^{n-2}(z)\right\}{\rm d}z.$
Using the notation of $F_{Y_{z}}$ in (3.1), the above identity becomes
$\mathbb{E}\left[\widehat{G}\right]={n}\int_{0}^{\infty}\left\{\int_{0}^{\infty}F_{Y_{z}}(t)\left[1-F_{Y_{z}}(t)\right]{\rm d}t\,\mathscr{L}_{X}^{n}(z)\right\}{\rm d}z.$ (3.2)
Since, for each fixed $z>0$, the mapping $t\in(0,\infty)\mapsto F_{Y_{z}}(t)$ in (3.1) is the CDF of $Y_{z}$, Proposition 2.1 yields
$\int_{0}^{\infty}F_{Y_{z}}(t)\left[1-F_{Y_{z}}(t)\right]{\rm d}t=\mathbb{E}[Y_{z}]G(Y_{z}),$ (3.3)
where $G(Y_{z})$ denotes the Gini coefficient of $Y_{z}$.
Finally, combining (3.2) and (3.3) completes the proof of the theorem. ∎
4 Characterization of the gamma family
The next result presents a characterization of the gamma family which, to the best of our knowledge, has not previously appeared in the literature.
Theorem 4.1.
Let $X$ be a nonnegative, absolutely continuous random variable belonging to a scale family, with density $f_{X}$. Assume that for every $z>0$ there exists a function $\xi(z)>0$ such that, for almost every $t>0$,
$\exp\{-zt\}f_{X}(t)=\frac{\mathscr{L}_{X}(z)}{\xi(z)}\,f_{X}\left(\frac{t}{\xi(z)}\right),$ (4.1)
where $\mathscr{L}_{X}(z)=\mathbb{E}[\exp\{-zX\}]$ denotes the Laplace transform of $X$, and $F_{X}$ is its corresponding CDF. Then $X\sim\text{Gamma}(\alpha,\lambda)$, for some $\alpha>0$, $\lambda>0$ (gamma distribution). More precisely,
$f_{X}(x)={\lambda^{\alpha}\over\Gamma(\alpha)}\,x^{\alpha-1}\exp\{-\lambda x\},\quad x>0,$
and necessarily
$\xi(z)=\frac{\lambda}{\lambda+z}.$
Conversely, if $X\sim\text{Gamma}(\alpha,\lambda)$, then identity (4.1) holds for all $z>0$.
Proof.
Since $X$ belongs to a scale family, there exists a baseline density $g$ such that
$f_{X}(x)=\frac{1}{\theta}\,g\left(\frac{x}{\theta}\right),$ (4.2)
for some $\theta>0$. Without loss of generality, assume that $\theta=1$. Equation (4.1) can therefore be rewritten as
$\exp\{-zt\}\,g\left({t}\right)=\frac{\mathscr{L}_{X}(z)}{\xi(z)}\,g\left(\frac{t}{\xi(z)}\right).$
Taking logarithms and differentiating with respect to $t$ yields
$-z+\left[\log(g({t}))\right]^{\prime}=\frac{1}{\xi(z)}\left[\log\left(g\left(\frac{t}{\xi(z)}\right)\right)\right]^{\prime}.$
Differentiating once more with respect to $t$, we obtain
$\left[\log(g\left({t}\right))\right]^{\prime\prime}=\frac{1}{\xi^{2}(z)}\left[\log\left(g\left(\frac{t}{\xi(z)}\right)\right)\right]^{\prime\prime}.$
Letting $x=t$ and $y=t/\xi(z)$, the above equation becomes
$x^{2}\bigl[\log g(x)\bigr]^{\prime\prime}=y^{2}\bigl[\log g(y)\bigr]^{\prime\prime},$
for all $x,y>0$. Hence, the quantity $x^{2}[\log g(x)]^{\prime\prime}$ must be constant. That is, there exists $C\in\mathbb{R}$ such that
$x^{2}\bigl[\log g(x)\bigr]^{\prime\prime}=C.$
Solving this differential equation yields
$g(x)=\exp\{K\}\,x^{-C}\exp\{Dx\},$ (4.3)
for some constants $K$, $C$, and $D$. Since $g$ is a probability density on $(0,\infty)$, it follows that C>1-C>-1 and $D<0$. Setting $\alpha=1-C$ and $\lambda=-D$, normalization together with (4.2) (with $\theta=1$) and (4.3) implies that $X\sim\mathrm{Gamma}(\alpha,\lambda)$.
The proof of the reciprocal implication is immediate and therefore omitted. ∎
Corollary 4.2.
Under the assumptions of Theorem 4.1,
$\frac{\mathbb{E}\left[\exp\{-zX\}\mathds{1}_{\{X\leqslant t\}}\right]}{\mathscr{L}_{X}(z)}=F_{X}\left(\frac{t}{\xi(z)}\right),\quad t>0;\qquad\xi(z)=\frac{\lambda}{\lambda+z},\quad z>0.$
5 Unbiasedness of the Gini estimator
Let $X$ be a nonnegative, absolutely continuous random variable belonging to a scale family, with density $f_{X}$, satisfying (4.1). Using the notation $F_{Y_{z}}$ introduced in (3.1), Theorem 4.1 implies that $X\sim\text{Gamma}(\alpha,\lambda)$ and Corollary 4.2 yields
$F_{Y_{z}}(t)=F_{X}\left(\frac{t}{\xi(z)}\right)=F_{\xi(z)X}(t),\quad\xi(z)=\frac{\lambda}{\lambda+z}.$
In other words,
$Y_{z}\stackrel{{\scriptstyle d}}{{=}}\xi(z)X\sim\text{Gamma}(\alpha,\lambda+z).$
Hence, $E[Y_{z}]=\alpha/(\lambda+z)$, $\mathscr{L}_{X}(z)=\lambda^{\alpha}/(\lambda+z)^{\alpha}$, and since the Gini coefficient is scale invariant, $G(Y_{z})=G(X)=G$. Applying Theorem 3.1, we obtain
$\mathbb{E}\left[\widehat{G}\right]=n\int_{0}^{\infty}\mathbb{E}[Y_{z}]\,G(Y_{z})\,\mathscr{L}_{X}^{\,n}(z)\,\mathrm{d}z=G\left[n\alpha\lambda^{n\alpha}\int_{0}^{\infty}{1\over(\lambda+z)^{n\alpha+1}}\,\mathrm{d}z\right]=G.$
That is, under gamma populations the Gini estimator $\widehat{G}$ is unbiased for the Gini coefficient $G$, as previously shown in Baydil et al., (2025) and Vila and Saulo, 2025a ; Vila and Saulo, 2025b .
6 Concluding remarks
In this paper, we developed a general framework for studying the expectation of the Gini coefficient estimator through Laplace-transform techniques. This framework led naturally to a characterization of the gamma family within the class of nonnegative scale families, based on the invariance of exponentially tilted distributions under scaling.
As a direct application, we provided an alternative proof of the unbiasedness of the Gini estimator under gamma populations. Compared with existing proofs, the proposed approach is conceptually simpler and emphasizes structural properties, namely, scale invariance and stability under exponential tilting, rather than distribution-specific algebraic calculations.
Beyond the gamma case, the representation derived in Theorem 3.1 may prove useful for studying bias properties of Gini-type estimators under other distributional settings, including mixtures and generalized scale families. Future research could explore whether similar characterizations arise for other inequality measures or whether approximate unbiasedness results can be obtained for broader classes of distributions.
Overall, the results reinforce the central role of the gamma family in inequality measurement and provide further insight into the probabilistic foundations of Gini estimation.
Acknowledgments
The research was supported in part by CNPq and CAPES grants from the Brazilian government.
Declaration
There are no conflicts of interest to disclose.